# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ Discussion
This study presents a state-space kinetic Ising model for estimating nonstationary and nonequilibrium neural dynamics, alongside a mean-field method for entropy flow estimation. Through analysis of mouse V1 neurons, we identified distinct field and coupling distributions across behavioral conditions. These structural shifts influenced entropy flow compositions in V1 neurons, revealing correlations with behavioral performance.
@@
To our knowledge, no inference methods have been proposed for time-dependent kinetic Ising models within the sequential Bayesian framework, which estimates parameters with uncertainty using optimized smoothness hyper-parameters (see [64] for Bayesian approach in a stationary case). While parameter estimation has often been considered under time-dependent fields with fixed couplings [40, 42] (see also [65, 66] for the equilibrium case), exceptions exist [41] that provide point estimates for time-varying couplings. These methods rely on mean-field equations relating equal-time and delayed correlations to coupling parameters, but estimating correlations at each time step is often infeasible in neuroscience data due to limited trial numbers in animal studies. Campajola et al. [67] proposed a point estimate of time-varying couplings using a score-driven method under the maximum likelihood principle but assumed all fields and couplings were uniformly scaled by a single time-varying parameter. In contrast, our state-space framework accommodates heterogeneous parameter dynamics and employs sequential Bayesian estimation with optimized smoothness parameters. These innovations are crucial for uncovering parameter variability’s impact on causal population dynamics and elucidating individual neuron’s contributions. In addition, our framework could be extended to analyze longer-term processes such as learning by treating time bins as trials within sessions and allowing parameters to vary across sessions, under the assumption of stationarity within each session. This would enable tracing learning trajectories of couplings among individual neurons when stable longitudinal recordings are available, an increasingly feasible scenario with recent advances in calcium imaging and electrophysiology[68, 69].
@@
Lower spike rates of V1 neurons observed during the active condition (see also [60]) contrast starkly with previous reports showing increased firing rates during active task engagement [70] or locomotion [71, 72]. Nevertheless, the diminished spike-rates found in the active condition (Fig. 6A and S1 Fig, S2 Fig) are in agreement with sparse population activity in processing natural images in mice V1 neurons [73, 74]. Further, active engagement broadened distributions of field and coupling parameters, possibly reflecting stronger and more diverse inputs from hidden neurons [75, 76]. These findings align with previously reported increased heterogeneous activities during the active condition and their correlation to behavioral performance [77]. The observed shift in cortical activity largely aligns with the effects of neuromodulators, such as acetylcholine (ACh)[78, 79] and norepinephrine (NE)[80], that alter local circuit interactions and global activity patterns, thereby regulating transitions such as quiet–active, and inattentive–attentive states [81, 82]. For example, Runfeldt et al.[78] demonstrated that spontaneous network events became sparser under ACh, as the probability of individual neurons participating in circuit activity was markedly reduced. In addition, ACh altered the temporal recruitment of neurons, delaying their activation relative to thalamic input and prolonging the window during which stereotyped activity propagated through local circuits. These findings indicate that ACh reorganizes cortical circuits into sparser and temporally extended modes of activity, potentially underlying the sparser population activity observed during task engagement and the stronger shift in temporal asymmetry in competent mice. However, we did not observe the previously reported decoupling of neuronal activity during active engagement (Fig. S5 Fig), which may suggest the involvement of additional mechanisms beyond those described above.
@@
In our analysis, the shift toward sparser activity during active engagement was significantly correlated with behavioral performance (Fig. 10B), consistent with sparse‑coding theories that posit efficient representations using few active neurons for natural images [83, 84, 85]. Moreover, mice with higher task performance exhibited greater entropy flow per spike during active compared with passive conditions (Fig. 10E), indicating that the capacity to form economical image representations via time‑asymmetric causal activity is also linked to behavioral performance. Future work should determine whether this pattern reflects a direct computational mechanism or a secondary consequence of network state (e.g., attention or arousal). Importantly, the proposed method further yields testable predictions for information coding. For instance, if entropy flow per spike indeed relates to computation, then (i) neurons whose receptive fields match the presented image features should show selectively higher entropy flow per spike, or (ii) population decoding accuracy should improve when restricting analysis to neurons with higher entropy flow per spike. These predictions provide avenues to experimentally validate the computational role of entropy flow.
@@
EEG, fMRI, and ECoG studies suggest that steady-state entropy production and related irreversibility metrics covary with consciousness level and cognitive load, and they reveal large-scale directed temporal structure [28, 86, 29, 87, 30, 88, 31, 89]. For example, in human fMRI, violations of the fluctuation–dissipation theorem are larger during wakefulness than deep sleep, and larger during tasks than rest [87]. Arrow-of-time analyses likewise show stronger temporal asymmetry during tasks than rest and identify a cortical hierarchy of asymmetry [88]. Our state-space kinetic Ising model complements these steady-state, macroscopic approaches by estimating entropy flow directly from spiking data without assuming stationarity, potentially illuminating the lower-level mechanisms of mesoscopic/microscopic circuit dynamics. In parallel, equilibrium Ising and energy-landscape methods have been successfully applied to binarized neuroimaging and electrophysiological signals to characterize correlation structure and attractor basins of large-scale brain networks [90, 91, 92, 93]. By contrast, our framework explicitly quantifies time-asymmetric entropy flow in non-stationary binary signals, complementing energy-landscape analyses of macroscopic stability with measures of time-dependent causal dynamics. In principle, the same approach could be extended upward in scale to local field potentials (LFPs), multi-electrode arrays (MEAs), or coarse-grained EEG/ECoG recordings, enabling multiscale analysis of nonequilibrium dynamics from circuit to whole-brain levels. However, the state-space method still faces limits in computational time and scale, constraining its use for these large-scale signals. Future improvements through parallelization, optimized algorithms, and refined mean-field approaches could extend its applicability and enhance entropy flow estimation.
@@
The kinetic Ising–based framework should also be viewed in light of its limitations. While analytically tractable, it imposes strong assumptions, pairwise couplings and conditional independence, that simplify neural dynamics but restrict interpretability. Our model-miss-specification analysis (Fig. 9) showed that reproducing the heavy-tailed spike-count statistics observed in real populations requires higher-order interactions; neglecting these leads to systematic biases, particularly in the tails. Likewise, synchronous updates imposed by the conditional independence obscure cascade-like recruitment within bins, producing bin-size–dependent distortions: large bins exaggerate heavy tails by merging cascades, while small bins preserve cascades but miss slower interactions distant in time. Thus, bin size entails a fundamental trade-off between resolving fine-scale recruitment and capturing broader dependencies.
@@
These limitations motivate extensions beyond the synchronous pairwise framework. The generalized linear models (GLMs) and related point-process models provide a natural asynchronous alternative, since spikes are modeled in fine-grained bins or continuous time and influence others through coupling kernels. However, entropy flow in such history-dependent systems requires full path probabilities, making estimation challenging. More broadly, fitted couplings and entropy flow should be regarded as statistical summaries of nonequilibrium dynamics, not direct measures of synaptic connectivity or mechanism. Future work must relax these constraints –by permitting asynchronous updates, incorporating higher-order dependencies, and developing principled estimators of entropy flow in non-Markovian settings –while remaining clear about the limits of inference when bridging statistical abstractions with physiology. For example, the alternating-shrinking higher-order interaction model (Eq. 66) could be extended to include asymmetric couplings, potentially with asynchronous updates in continuous-time limit.
@@
In summary, by developing a state-space kinetic Ising model that accounts for both non-stationary and nonequilibrium properties, we have demonstrated how task engagement modulates neuronal firing activity and coupling diversity. Our approach incorporates time-varying entropy flow estimation, revealing that time-asymmetric, irreversible activity emerges within sparsely active populations during task engagement—an effect correlated with the mouse’s behavioral performance. These findings underscore the utility of our approach, offering new insights into the thermodynamic underpinnings of neural computation.
@@ Methods
We summarize the expectation-maximization algorithm for estimating the state-space kinetic Ising model with optimized hyper-parameters. See Supporting Information 1 for their derivations.
E-step Given the optimal parameter $\mathbf{w}$, we obtain the estimate of the state $\boldsymbol{\theta}$ given all the data available. When estimating the parameter $\boldsymbol{\theta}^{i}_{t}$ ($t=0,1,\ldots,T$, $i=1,\ldots,N$ ) from the spike data $\mathbf{x}_{t}$ ($t=0,1,\ldots,T$), we first obtain the filter distribution by the sequentially applying the Bayes theorem:
$p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})$ $=\frac{p(\mathbf{x}_{t}|\boldsymbol{\theta}_{t},\mathbf{x}_{0:t-1},\mathbf{w})p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})}{p(\mathbf{x}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})}$ (22)
Here the one-step prediction density is computed using the Chapman-Kolmogorov equation:
$p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})=\prod_{i=1}^{N}\int p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1},\mathbf{Q}^{i})p(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{t-1})d\boldsymbol{\theta}^{i}_{t-1}$ (23)
By assuming that the filter density at the previous time step t1t-1 is given by the Gaussian density with mean $\boldsymbol{\theta}^{i}_{t-1|t-1}$ and covariance $\mathbf{W}^{i}_{t-1|t-1}$. The one-step prediction density becomes the Gaussian density given by
$p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})$ $=\prod_{i=1}^{N}\exp\left[-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}\mathbf{(}\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t|t-1})\right],$ (24)
where the mean and covariance are given by $\boldsymbol{\theta}^{i}_{t|t-1}=\boldsymbol{\theta}^{i}_{t-1|t-1}$ and $\mathbf{W}^{i}_{t|t-1}=\mathbf{W}^{i}_{t-1|t-1}+\mathbf{Q}^{i}$. Then, the filter density is given by
@@
$\phantom{==}\times\prod_{i=1}^{N}\exp\left[-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})\right].$ (25)
@@
$\boldsymbol{\theta}^{i}_{t|t}=\arg\max_{\boldsymbol{\theta}^{i}_{t}}\log p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t},\mathbf{w})$ (26)
for $i=1,\ldots,N$, while the variance is approximated using the Hessian:
$\mathbf{W}^{i}_{t|t}$ $=\left[-\frac{\partial}{\partial\boldsymbol{\theta}_{t}^{i}}\frac{\partial\log p(\boldsymbol{\theta}_{t}^{i}|\mathbf{x}_{0:t},\mathbf{w})}{\partial\left(\boldsymbol{\theta}_{t}^{i}\right)^{\mathsf{T}}}\bigg|_{\boldsymbol{\theta}_{t}^{i}=\boldsymbol{\theta}^{i}_{t|t}}\right]^{-1}$
@@
where $\mathbf{G}(\boldsymbol{\theta}^{i}_{t|t})\equiv\left.\frac{\partial^{2}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}_{t}^{i}\partial\left(\boldsymbol{\theta}_{t}^{i}\right)^{\mathsf{T}}}\right|_{\hat{\theta}^{i}_{t}=\theta^{i}_{t|t}}$ is the Fisher information matrix with respect to $\boldsymbol{\theta}^{i}_{t}$ computed for the kinetic Ising model over the trials. We searched the MAP estimate by the Newton-Raphson method utilizing the Hessian evaluated at a search point.
Next, we obtain the smoother density by recursively applying the formula. Because the filter density and state transitions are approximated by normal distributions, we follow the fixed-interval smoothing algorithm developed for the Gaussian distributions [94]. In this method, the smoothed mean and covariance are recursively obtained by the following equations:
@@
$\mathbf{A}^{i}_{t-1}$ $=\mathbf{W}^{i}_{t-1|t-1}\left(\mathbf{W}^{i}_{t|t-1}\right)^{-1}$ (30)
@@
M-step When optimizing the hyperparameter $\mathbf{Q}^{i}$ in M-step, the update formula used is obtained as follows
$\mathbf{Q}^{i}=\frac{1}{T-1}$ $\sum_{t=2}^{T}\left((\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|T})(\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|T})^{\top}\right$.
$\quad\left.-\boldsymbol{\theta}^{i}_{t|T}\boldsymbol{\theta}^{i\top}_{t-1|T}-\mathbf{W}^{i}_{t,t-1|T}+\boldsymbol{\theta}^{i}_{t-1|T}\boldsymbol{\theta}^{i\top}_{t-1|T}+\mathbf{W}^{i}_{t-1|T}\right).$ (31)
We compute the lag-one smoothing covariance matrix ${W}^{i}_{t,t-1|T}$ following the method of De Jong and Mackinnon [95], $\mathbf{W}^{i}_{t,t-1|T}=\mathbf{W}^{i}_{t|t}(\mathbf{W}^{i}_{t+1|t})^{-1}\mathbf{W}^{i}_{t|T}$. We also note that the optimization of a diagonal of the form $\mathbf{Q}^{i}={\rm diag}[\lambda^{i}_{0},\ldots,\lambda^{i}_{N}]$ or $\mathbf{Q}^{i}=\lambda^{i}\mathbf{I}$ can be performed by taking diagonal and trace of the r.h.s of the equation above, respectively.
@@
The convergence of the EM algorithm is assessed by computing the approximate log-likelihood function (Eq.6) using the Laplace approximation. Using Eqs. 1 and 24 in Eq.6, the approximate log-likelihood function for the hyperparameter $\mathbf{w}$ is obtained as
$\log[p(\mathbf{x}_{0:T}|\mathbf{w})]=\log[p(\mathbf{x}_{0}|\mathbf{w})]+\sum_{t=2}^{T}\sum_{i=1}^{N}[\frac{1}{2}\log|\mathbf{W}^{i}_{t|t}|-\frac{1}{2}\log|\mathbf{W}^{i}_{t|t-1}|+q(\boldsymbol{\theta}^{i}_{t|t})].$ (33)
@@
Here we extend the mean-field approximation method developed for the steady-state kinetic Ising model [43] to make it applicable to non-stationarity systems.
@@
We calculate these conditional entropies using the Gaussian approximation as follows. First, ${\sigma}_{t}^{\rm forward}$ can represented as
${\sigma}_{t}^{\rm forward}$ $=-\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}\sum_{\mathbf{x}_{t-2}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})p(\mathbf{x}_{t-2})\log p(\mathbf{x}_{t}|\mathbf{x}_{t-1})$
$=-\sum_{\mathbf{x}_{t-1}}\sum_{\mathbf{x}_{t-2}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})p(\mathbf{x}_{t-2})\sum_{\mathbf{x}_{t}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})\log p(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ (37)
Here, we replace $p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})$ with an independent model $Q(\mathbf{x}_{t-1})$ defined as
$Q(\mathbf{x}_{t-1})=\prod_{i}Q(x_{i,t-1})=\prod_{i}\frac{e^{x_{i,t}\theta_{i,t-1}^{0}}}{1+e^{\theta_{i,t-1}^{0}}},$ (38)
where $\theta_{i,t-1}^{0}$ is the parameter of the independent model for the $i$-th neuron. Then, ${\sigma}_{t}^{\rm forward}$ is approximated as
${\sigma}_{t}^{\rm forward}$ $\simeq-\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})\sum_{\mathbf{x}_{t}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})\log p(\mathbf{x}_{t}|\mathbf{x}_{t-1}).$ (39)
@@
$p(x_{i,t}|\mathbf{x}_{t-1})=\frac{e^{x_{i,t}h_{i,t}(\mathbf{x}_{t-1})}}{1+e^{h_{i,t}(\mathbf{x}_{t-1})}}$ (41)
@@
$p(x_{i,t}$ $=1|\mathbf{x}_{t-1})=\frac{e^{h_{i,t}(\mathbf{x}_{t-1})}}{1+e^{h_{i,t}(\mathbf{x}_{t-1})}}=r(h_{i,t}(\mathbf{x}_{t-1})),$ (44)
$p(x_{i,t}$ $=0|\mathbf{x}_{t-1})=\frac{1}{1+e^{h_{i,t}(\mathbf{x}_{t-1})}}=1-r(h_{i,t}(\mathbf{x}_{t-1})).$ (45)

### v3

@@ Discussion
This study presents a state-space kinetic Ising model for estimating nonstationary and nonequilibrium neural dynamics and introduces a mean-field method for entropy flow estimation. Through analysis of mouse V1 neurons, we identified distinct field and coupling distributions across behavioral conditions. These structural shifts influenced entropy flow compositions in V1 neurons, revealing correlations with behavioral performance.
@@
To our knowledge, no inference methods have been proposed for time-dependent kinetic Ising models within the sequential Bayesian framework, which estimates parameters with uncertainty using optimized smoothness hyperparameters (see [64] for a Bayesian approach in a stationary case). While parameter estimation has often been considered under time-dependent fields with fixed couplings [40, 42] (see also [65, 66] for the equilibrium case), exceptions exist [41] that provide point estimates for time-varying couplings. These methods rely on mean-field equations relating equal-time and delayed correlations to coupling parameters, but estimating correlations at each time step is often infeasible in neuroscience data due to limited trial numbers in animal studies. Campajola et al. [67] proposed a point estimate of time-varying couplings using a score-driven method under the maximum likelihood principle, but assumed all fields and couplings were uniformly scaled by a single time-varying parameter. In contrast, our state-space framework accommodates heterogeneous parameter dynamics and employs sequential Bayesian estimation with optimized smoothness parameters. These innovations are crucial for uncovering parameter variability’s impact on causal population dynamics and elucidating individual neurons’ contributions.
@@
Lower spike rates of V1 neurons observed during the active condition (see also [60]) contrast starkly with previous reports showing increased firing rates during active task engagement [68] or locomotion [69, 70]. Nevertheless, the diminished spike-rates found in the active condition (Fig. 6A and Supplementary Fig. S1, S2) are in agreement with sparse population activity in processing natural images in mouse V1 neurons [71, 72]. Further, active engagement broadened distributions of field and coupling parameters, possibly reflecting stronger and more diverse inputs from hidden neurons [73, 74]. These findings align with previously reported increased heterogeneous activities during the active condition and their correlation to behavioral performance [75]. The observed shift in cortical activity largely aligns with the effects of neuromodulators, such as acetylcholine (ACh) [76, 77] and norepinephrine (NE) [78], that alter local circuit interactions and global activity patterns, thereby regulating transitions such as quiet-active, and inattentive-attentive states [79, 80]. For example, Runfeldt et al. [76] demonstrated that spontaneous network events became sparser under ACh, as the probability of individual neurons participating in circuit activity was markedly reduced. In addition, ACh altered the temporal recruitment of neurons, delaying their activation relative to thalamic input and prolonging the window during which stereotyped activity propagated through local circuits. These findings indicate that ACh reorganizes cortical circuits into sparser and temporally extended modes of activity, potentially underlying the sparser population activity observed during task engagement and the stronger shift in entropy flow per spike in competent mice. However, we did not observe the previously reported decoupling of neuronal activity during active engagement (Supplementary Fig. S5), which may suggest the involvement of additional mechanisms beyond those described above.
@@
In our analysis, the shift toward sparser activity during active engagement was significantly correlated with behavioral performance (Fig. 10B), consistent with sparse-coding theories that posit efficient representations using a few active neurons for natural images [81, 82, 83]. Moreover, mice with higher task performance exhibited greater entropy flow per spike during active compared with passive conditions (Fig. 10E), indicating that the capacity to form economical image representations via time-asymmetric causal activity is also linked to behavioral performance. Future work should determine whether this pattern reflects a direct computational mechanism or a secondary consequence of network state (e.g., attention or arousal). Importantly, the proposed method further yields testable predictions for information coding. For instance, if entropy flow per spike indeed relates to computation, then (i) neurons whose receptive fields match the presented image features should show selectively higher entropy flow per spike, or (ii) population decoding accuracy is expected to remain largely unchanged when the analysis is restricted to neurons with higher entropy flow per spike. Moreover, targeted pharmacological or optogenetic manipulations of neuromodulatory systems are predicted to induce systematic changes in entropy flow by modulating coupling variability, thereby altering coding efficiency. These predictions provide avenues to experimentally validate the computational role of entropy flow.
@@
EEG, fMRI, and ECoG studies suggest that steady-state entropy production and related irreversibility metrics covary with consciousness level and cognitive load, and they reveal large-scale directed temporal structure [28, 84, 29, 85, 30, 86, 31, 87]. For example, in human fMRI, violations of the fluctuation-dissipation theorem are larger during wakefulness than deep sleep, and larger during tasks than rest [85]. Arrow-of-time analyses likewise show stronger temporal asymmetry during tasks than rest and identify a cortical hierarchy of asymmetry [86]. Our state-space kinetic Ising model complements these steady-state, macroscopic approaches by estimating entropy flow directly from spiking data without assuming stationarity, potentially illuminating the lower-level mechanisms of mesoscopic/microscopic circuit dynamics. In parallel, equilibrium Ising and energy-landscape methods have been successfully applied to binarized neuroimaging and electrophysiological signals to characterize correlation structure and attractor basins of large-scale brain networks [88, 89, 90, 91]. Our framework explicitly quantifies time-asymmetric entropy flow in nonstationary binary signals, complementing energy-landscape analyses of macroscopic stability with measures of time-dependent causal dynamics. In principle, our approach could be extended upward in scale to local field potentials (LFPs), multi-electrode arrays (MEAs), or coarse-grained EEG/ECoG recordings, enabling multiscale analysis of nonequilibrium dynamics from circuit to whole-brain levels.
In addition, our framework could be extended to analyze longer-term processes such as learning by treating time bins as trials within sessions and allowing parameters to vary across sessions, under the assumption of stationarity within each session. This would enable tracing learning trajectories of couplings among individual neurons when stable longitudinal recordings are available, an increasingly feasible scenario with recent advances in calcium imaging and electrophysiology [92, 93]. However, the state-space method still faces limits in computational time and scale, constraining its use for large-scale signals. Future improvements through parallelization, optimized algorithms, and refined mean-field approaches could extend its applicability and enhance entropy flow estimation.
@@
The kinetic Ising-based framework should also be viewed in light of its theoretical limitations. While analytically tractable, it imposes strong assumptions – namely, pairwise couplings and conditional independence – that simplify neural dynamics but restrict interpretability. Our model misspecification analysis (Fig. 5) showed that reproducing the heavy-tailed spike-count statistics observed in real populations requires higher-order interactions; neglecting these leads to systematic biases, particularly in the tails. Likewise, synchronous updates imposed by conditional independence obscure cascade-like recruitment within bins in experimental data, leading to bin-size-dependent distortions: Large bins capture heavy tails by merging cascades, which the model fails to represent, while small bins preserve fine-scale cascades, but the model misses slower interactions distant in time. These limitations motivate extensions beyond the synchronous pairwise framework. The generalized linear models (GLMs) and related point-process models provide a natural asynchronous alternative with longer history-dependency, since spikes are modeled in fine-grained bins or continuous time and influence others through coupling kernels. However, entropy flow in such history-dependent systems requires full path probabilities, making estimation challenging.
@@
More broadly, fitted couplings and entropy flow should be regarded as statistical summaries of nonequilibrium dynamics, not direct measures of synaptic connectivity or mechanism. Future work must relax these constraints –by permitting asynchronous updates, incorporating higher-order dependencies, and developing principled estimators of entropy flow in non-Markovian settings –while remaining clear about the limits of inference when bridging statistical abstractions with physiology. For example, the alternating-shrinking higher-order interaction model (Eq. 63) could be extended to include asymmetric couplings, potentially with asynchronous updates in a continuous-time limit.
@@
In summary, by developing a state-space kinetic Ising model that accounts for both nonstationary and nonequilibrium properties, we have demonstrated how task engagement modulates neuronal firing activity and coupling diversity. Our approach incorporates time-varying entropy flow estimation, revealing that time-asymmetric, irreversible activity emerges within sparsely active populations during task engagement—an effect correlated with the mouse’s behavioral performance. These findings underscore the utility of our approach, offering new insights into the thermodynamic underpinnings of neural computation.
@@ Methods
We summarize the expectation-maximization algorithm for estimating the state-space kinetic Ising model with optimized hyperparameters. See Supplementary Note 1 for more details.
E-step: Given the hyperparameters $\mathbf{w}$, we obtain the estimate of the state $\boldsymbol{\theta}_{t}$ given all the data available. When estimating the parameters $\boldsymbol{\theta}^{i}_{t}$ ($t=0,1,\ldots,T$, $i=1,\ldots,N$ ) from the spike data $\mathbf{x}_{t}$ ($t=0,1,\ldots,T$), we first obtain the filter density by the sequentially applying the Bayes theorem:
$p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})$ $=\frac{p(\mathbf{x}_{t}|\boldsymbol{\theta}_{t},\mathbf{x}_{0:t-1},\mathbf{w})p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})}{p(\mathbf{x}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})}.$ (22)
Here, the one-step prediction density is computed using the Chapman-Kolmogorov equation:
$p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})=\prod_{i=1}^{N}\int p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1},\mathbf{Q}^{i})p(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{t-1})d\boldsymbol{\theta}^{i}_{t-1}.$ (23)
By assuming that the filter density for the $i$-th neuron at the previous time step t1t-1 is given by the Gaussian distribution with mean $\boldsymbol{\theta}^{i}_{t-1|t-1}$ and covariance $\mathbf{W}^{i}_{t-1|t-1}$, the one-step prediction density becomes the Gaussian distribution whose mean $\boldsymbol{\theta}^{i}_{t|t-1}$ and covariance $\mathbf{W}^{i}_{t|t-1}$ are given by
$\boldsymbol{\theta}^{i}_{t|t-1}$ $=\boldsymbol{\theta}^{i}_{t-1|t-1},$ (24)
$\mathbf{W}^{i}_{t|t-1}$ $=\mathbf{W}^{i}_{t-1|t-1}+\mathbf{Q}^{i},$ (25)
with $\boldsymbol{\theta}^{i}_{1|0}=\boldsymbol{\mu}^{i}$ and $\mathbf{W}^{i}_{1|0}=\boldsymbol{\Sigma}^{i}$ being the hyperparameters of the initial Gaussian distribution, $p(\boldsymbol{\theta}_{1}^{i}|\boldsymbol{\mu}^{i},\boldsymbol{\Sigma}^{i})$. Then, the filter density is given as
@@
$\phantom{=}\cdot\prod_{i=1}^{N}\exp\left[-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})\right].$ (26)
@@
$\boldsymbol{\theta}^{i}_{t|t}=\arg\max_{\boldsymbol{\theta}^{i}_{t}}\log p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t},\mathbf{w}),$ (27)
for $i=1,\ldots,N$, while the covariance is approximated using the Hessian as
$\mathbf{W}^{i}_{t|t}$ $=\left[-\frac{\partial^{2}\log p(\boldsymbol{\theta}_{t}^{i}|\mathbf{x}_{0:t},\mathbf{w})}{\partial\boldsymbol{\theta}_{t}^{i}\partial\left(\boldsymbol{\theta}_{t}^{i}\right)^{\mathsf{T}}}\bigg|_{\boldsymbol{\theta}_{t}^{i}=\boldsymbol{\theta}^{i}_{t|t}}\right]^{-1}$
@@
where $\mathbf{G}(\boldsymbol{\theta}^{i}_{t})\equiv\sum_{l=1}^{L}\left.\frac{\partial^{2}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}_{t}^{i}\partial\left(\boldsymbol{\theta}_{t}^{i}\right)^{\mathsf{T}}}\right|_{\boldsymbol{\theta}^{i}_{t}=\boldsymbol{\theta}^{i}_{t|t}}$ is the Fisher information matrix with respect to $\boldsymbol{\theta}^{i}_{t}$ computed for the kinetic Ising model over the trials. We computed the MAP estimate by the Newton-Raphson method utilizing the Hessian evaluated at a search point.
Next, we obtain the smoother density by recursively applying the formula below. Because the filter density and state transitions are approximated by normal distributions, we follow the fixed-interval smoothing algorithm developed for the Gaussian distributions [94]. In this method, the smoothed mean and covariance are recursively obtained by the following equations:
@@
$\mathbf{A}^{i}_{t-1}$ $=\mathbf{W}^{i}_{t-1|t-1}\left(\mathbf{W}^{i}_{t|t-1}\right)^{-1},$ (31)
@@
M-step: We optimize the hyperparameters given the smoothed posteriors. To optimize the hyperparameter $\mathbf{Q}^{i}$, we used the following update formula that maximizes the lower bound of the log marginal likelihood:
$\mathbf{Q}^{i}$ $=\frac{1}{T-1}\sum_{t=2}^{T}\left[(\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|T})(\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|T})^{\top}\right$.
$\phantom{=}\left.+\mathbf{W}^{i}_{t|T}-\mathbf{W}^{i}_{t-1,t|T}-\mathbf{W}^{i}_{t,t-1|T}+\mathbf{W}^{i}_{t-1|T}\right].$ (32)
We compute the lag-one smoothing covariance matrix $\mathbf{W}^{i}_{t,t-1|T}$ following the method of De Jong and Mackinnon [95]: $\mathbf{W}^{i}_{t,t-1|T}=\mathbf{W}^{i}_{t|t}(\mathbf{W}^{i}_{t+1|t})^{-1}\mathbf{W}^{i}_{t|T}$. We also note that the optimization of a diagonal of the form $\mathbf{Q}^{i}={\rm diag}[\lambda^{i}_{0},\ldots,\lambda^{i}_{N}]$ or $\mathbf{Q}^{i}=\lambda^{i}\mathbf{I}$ can be performed by taking diagonal and trace of the r.h.s of the equation above, respectively.
@@
The convergence of the EM algorithm is assessed by computing the approximate log marginal likelihood function (Eq.6) using the Laplace approximation. Using the mean and covariance of the filter and one-step prediction densities, the approximate log marginal likelihood function for the hyperparameters $\mathbf{w}$ is obtained as
$\log p(\mathbf{x}_{0:T}|\mathbf{w})=\log p(\mathbf{x}_{0})$
$+\sum_{t=1}^{T}\sum_{i=1}^{N}\left[\frac{1}{2}\log|\mathbf{W}^{i}_{t|t}|-\frac{1}{2}\log|\mathbf{W}^{i}_{t|t-1}|+q(\boldsymbol{\theta}^{i}_{t|t})\right].$ (34)
See Supplementary Note 1 for the derivations, and the functional form of $q(\cdot)$.
@@
Here we extend the mean-field approximation method developed for the steady-state kinetic Ising model [43] to make it applicable to nonstationary systems.
@@
We calculate these conditional entropies using the Gaussian approximation as follows.
We begin with approximating the forward conditional entropy as
${\sigma}_{t}^{\rm forward}$ $=-\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})p(\mathbf{x}_{t-1})\log p(\mathbf{x}_{t}|\mathbf{x}_{t-1})$
$\simeq-\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})\sum_{\mathbf{x}_{t}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})\log p(\mathbf{x}_{t}|\mathbf{x}_{t-1}).$ (38)
Here we replaced $p(\mathbf{x}_{t-1})$ with an independent model $Q(\mathbf{x}_{t-1})$ defined as
$Q(\mathbf{x}_{t-1})=\prod_{i}Q(x_{i,t-1}).$ (39)
@@
$p(x_{i,t}|\mathbf{x}_{t-1})$ $=e^{x_{i,t}h_{i,t}(\mathbf{x}_{t-1})-\psi(h_{i,t}(\mathbf{x}_{t}))}$ (41)
@@
Here, we redefined the log normalization function $\psi$ as a function of $h_{i,t}(\mathbf{x}_{t})$: $\psi(h_{i,t}(\mathbf{x}_{t}))=\log(1+e^{h_{i,t}(\mathbf{x}_{t})})$.
@@
$p(x_{i,t}$ $=1|\mathbf{x}_{t-1})=r(h_{i,t}(\mathbf{x}_{t-1})),$ (44)
$p(x_{i,t}$ $=0|\mathbf{x}_{t-1})=1-r(h_{i,t}(\mathbf{x}_{t-1})).$ (45)